Write each of the following as decimals:
3 + 7/10??????????

Grade:12

2 Answers

Harshit Singh
2 years ago
Dear Student

3 + 7 / 10
= 3 + 0.7
= 3.7

Thanks
Shraddha Singh
one year ago
Dear Student,
 
3 + 7/10
= 3/1 + 7/10
Taking LCM as 10,
= (3×10) + (7×1) / 10
= (30 + 7)/10
= 37/10
= 3.7
 
Hope it helps!
Thank you